# Nuclear Cold QCD: Review and Future Strategy

### v1

@@ IV Radiative parton energy loss
Reactions with nuclei are essential for the study of QCD. Cold nuclear matter effects can be investigated in $e+A$ [24] and $p+A$ collisions [369, 370]. The modification of hadron production cross sections at low and intermediate $p_{T}$, known as the Cronin effect [37], was established as early as the mid 1970s. In the 1990s, fixed-target measurements of Drell-Yan and charmonia at Fermilab (see Sec. II) found sizeable cross section suppression at large $x_{F}$. While initially attributed only to nPDF effects, it was noted early on that the attenuation of the production rate of such states is also compatible with CNM energy loss [371].
In the late 1990s and early 2000s important developments in many-body perturbative QCD produced theoretical frameworks to evaluate non-Abelian parton energy loss due to soft gluon emission [372, 373, 19, 10, 374, 375, 376, 377]. Subsequently, energy loss has been extensively employed as a mechanism for the modification of cross sections and other, related, observables in nuclear matter. This approach has been applied to both Drell-Yan [378, 41, 379, 42] and $J/\psi$ [371, 380, 13, 381] suppression at large $x_{F}$ in minimum bias $p+A$ collisions as well as jet modification [382] in central $p+A$ collisions at very high energies.
In the remainder of this section, we describe the various energy loss regimes and observables that are sensitive to them. We further touch upon recent progress on understanding in-medium radiation beyond the soft gluon emission limit and its connection to other perturbative approaches such as saturation.
@@ IV.1 Energy loss regimes
These processes are widely studied as a dominant mechanism behind high-$p_{T}$ hadron and jet suppression in heavy-ion collisions. However, they may also play a significant role in forward particle production of quarkonium [380, 13, 381], light hadrons [383, 384], and heavy flavor mesons [231] in $p+A$ collisions.
@@
The difference between gluon radiation in protons and nuclei, the medium-induced gluon radiation spectrum, exhibits different scaling properties depending on the typical gluon formation time $t_{f}\sim\omega/k_{T}^{2}$, where $\omega$ and $k_{T}$ are the energy and the transverse momentum of the radiated gluon, respectively. This gluon formation time needs to be compared to the two other length scales appearing in the process, the parton mean free path in the medium, $\lambda$, and the length of the medium traversed, $L$. In practice, qualitative analytic behavior can be derived in the limits of very large parton energy and large numbers of scatterings while numerical corrections can be large for lower energies and a smaller number of scatterings [10, 385]. Nevertheless, it is useful to recall the different regimes identified in the medium-induced gluon spectrum depending on the gluon formation time $t_{f}$ [386]:
  • (i)
@@
    In the incoherent or Bethe-Heitler regime, $t_{f}\lesssim\lambda$, each scattering center acts as an independent radiation source. This regime is at work for very soft gluons, $\omega\ll\mu^{2}\lambda=\hat{q}\lambda^{2}$, where $\hat{q}=\mu^{2}/\lambda$ is known as the medium transport coefficient and $\mu$ is the inverse of the interaction range [386, 387];
  • (ii)
@@
    In the partially coherent or Landau-Pomeranchuk-Migdal (LPM) regime, $\lambda\ll t_{f}\lesssim L$, a group of scattering centers with $t_{f}/\lambda$ acts as a single radiator, leading to relative suppression of the gluon radiation intensity with respect to the intensity in the incoherent regime. The above condition translates into $\mu^{2}\lambda\ll\omega\lesssim\hat{q}L^{2}$. The average LPM energy loss scales parametrically as $\Delta E_{{}_{\textnormal{LPM}}}\propto\hat{q}L^{2}\ln(E/\mu^{2}L)$, with an at-most logarithmic dependence on the parton energy;
  • (iii)
@@
    In the fully coherent regime, $t_{f}\gg L$, all scattering centers in the medium act as a coherent radiation source. In this large gluon energy regime, $\omega\gg\hat{q}L^{2}$, the average energy loss $\Delta E_{{}_{\textnormal{FCEL}}}\propto(\hat{q}\,L)^{1/2}\,(E/M_{T})$ becomes proportional to the parton energy $E$ [12, 13], overwhelmsing LPM energy loss, $\Delta E_{{}_{\textnormal{FCEL}}}\gg\Delta E_{{}_{\textnormal{LPM}}}$. The phenomenology of FCEL has been investigated in quarkonium [380, 13, 381], light hadron [383, 384], and heavy flavor meson [231] production in $p+A$ collisions. The effects of FCEL, which explicitly break QCD factorization, are sizable and therefore cannot be neglected. Given this, including measurements of hadron production in $p+A$ collisions in global analyses of nPDFs [223, 222, 224] without proper implementation of FCEL might lead to biased nPDF extractions.
  • (iv)
@@
    Last but not least, initial-state interactions can partially cancel radiation effects, especially at low frequencies. This type of medium-induced radiation [377] has Bertsch-Gunion (BG) [388, 389] scaling properties, $\omega(dI/d\omega)\sim L$, without additional power suppressions. The type of energy loss is manifested in Drell-Yan, hadron, and jet suppression at large $x_{F}$ [41, 390, 391].
@@
Figure 12: Schematic log-log plot showing the parametric dependence of the induced radiation spectrum $\omega\,(dI/d\omega)$ divided by $\alpha_{s}$, for parton production in the color representation $R$ at small angles in the regimes of gluons formation times described in this section. Taken from [392].
@@
A schematic plot summarizes the parametric dependence of the medium-induced radiation spectrum in Fig. 12. It can be shown that if a parton suddenly decelerated (accelerated), leading to initial-state (final-state) energy loss, the medium-induced gluon spectrum decreases as $\omega(dI/d\omega)\sim\omega^{-1}$ at $\omega\gg\hat{q}\,L^{2}$, leading to $\Delta E_{{}_{\textnormal{IS}}}=\Delta E_{{}_{\textnormal{FS}}}=\Delta E_{{}_{\textnormal{LPM}}}$. On the contrary, whenever the hard process appears to be the small-angle deflection of a high energy parton, as is typically the case for hadron production in $p+A$ collisions, when both the initial-state and the final-state partons carry a color charge (see Fig. 13), the associated medium-induced gluon radiation arises from the interference of gluon radiation before and after the hard scattering. In that case, the induced spectrum is dominated by the fully coherent radiation regime, $t_{\textnormal{f}}\gg L$, leading to $\Delta E=\Delta E_{{}_{\textnormal{FCEL}}}$.
@@
Figure 13: Induced gluon radiation from a fast parton experiencing multiple scatterings in a dense nucleus taken from [392].
@@
Disentangling the different regimes of radiative parton energy loss is key to understanding its effects. At high collision energies, typically at the LHC, the effect of LPM energy loss in nuclear matter should be negligible because the incoming parton energy $E$ is extremely large in the nuclear rest frame, leading to a negligible fractional energy loss, $\Delta E_{{}_{\rm LPM}}/E\sim\ln E/E\ll 1$. In this respect, the LPM regime would be best probed in low energy nuclear collisions.
@@
Initial state $\hat{q}\,L^{2}\,\ln E$ Drell-Yan $h+A$ E906, COMPASS [42]
Final state $\hat{q}\,L^{2}\,\ln E$ $h$, jets $e+A$ CLAS, HERMES, EIC [393]
Fully coherent $(\hat{q}\,L)^{1/2}\,(E/M_{T})$ $h$, jets $h+A$ SPS, FNAL, RHIC, LHC [380, 13, 381]
@@
Table 6: Energy loss processes in collisions on nuclear targets, their parametric dependencies, and the systems and facilities in which they might be probed.
@@
IV.2 The cold nuclear matter transport coefficient $\hat{q}$
The strength of induced gluon radiation in cold nuclear matter is governed by the transport coefficient $\hat{q}$ within the Baier–Dokshitzer–Mueller–Peigné–Schiff and Zakharov (BDMPS–Z) formalism [^7] Note that in (5), $\tilde{\mu}^{2}$ already accounts for the typical transverse momentum squared in two dimensions. [386, 19, 394, 387], which quantifies the scattering power of the nuclear medium. It is related to the gluon distribution $xG(x,Q^{2})$ inside each nucleon,
$\hat{q}\equiv\frac{\tilde{\mu}^{2}}{\lambda}=\frac{4\pi^{2}\,\alpha_{s}\,C_{\mathrm{R}}}{N_{c}^{2}-1}\;\rho\,xG\bigl{(}x,Q^{2}\bigr{)}\,,$ (5)
where $Q^{2}\sim\Delta p_{T}^{2}$ denotes the characteristic transverse momentum scale, $C_{\mathrm{R}}$ is the parton color factor —with $C_{\mathrm{F}}=(N_{c}^{2}-1)/2N_{c}$ for a quark and $C_{\mathrm{A}}=N_{c}$ for a gluon— and $\rho$ is the nuclear density.
@@ IV.2 The cold nuclear matter transport coefficient $\hat{q}$
The value of $\hat{q}$ can be determined from different processes, such as fully coherent energy loss, initial-state and final-state energy loss, and transverse momentum broadening. Uncertainties in its determination arise from the type of energy loss formalism used and how it is implemented phenomenologically as well as the treatment of the nuclear geometry. Quite remarkably, a number studies find roughly the same value of $\hat{q}$ within uncertainties. FCEL studies of $J/\psi$ suppression in $p+A$ collisions [13] determine a value of $\hat{q}_{0}\equiv\hat{q}(x=10^{-2})$ of $0.05\lesssim\hat{q}_{0}\lesssim 0.09\,\text{GeV}^{2}/\text{fm}$ depending on the nPDF effects. Using these values in the calculation of initial-state quark energy loss leads to a satisfactory description of the preliminary E906 Drell-Yan $p+A$ data [42]. The value of $\hat{q}_{0}$ can also be determined from studies of the transverse momentum broadening experienced by a particle produced in nuclear collisions, $\Delta p_{T}^{2}$, see e.g. Ref. [14],
@@
This value indirectly related to multiple parton scattering in the nuclear medium, $\Delta p_{T}^{2}\propto\hat{q}_{A}L_{A}-\hat{q}_{p}L_{p}$. As shown in Fig. 14, using world data on $J/\psi$ and Drell-Yan transverse momentum broadening in $h+A$ collisions from SPS to the LHC, the transport coefficient is found to be $\hat{q}_{0}=0.051\,\text{GeV}^{2}/\text{fm}$ or $\hat{q}_{0}=0.075\,\text{GeV}^{2}/\text{fm}$ depending on the assumed color state of the produced $c\overline{c}$ pair that fragments into a $J/\psi$ [15]. The lower value of $\hat{q}^{2}_{0}$ assumes color octet production while the larger value assumed the pair is produced as a color singlet.
@@
IV.3 A renormalization group approach to energy loss
In the past decades, successful energy loss phenomenology was developed for both hot and cold QCD matter. Such theoretical models, however, have clear limitations. They cannot be systematically improved to higher perturbative orders, and do not give a correct description of in-medium parton showers. For example, in the case of open heavy flavor production the $g\rightarrow Q\bar{Q}$ process, which is especially important at high energies, is not included. To address these deficiencies, full medium-induced splitting kernels [396, 397] were derived using effective field theory and lightcone wave function approaches. These analogues of DGLAP splitting functions in matter further require implementations in cross section evaluation that go beyond energy loss phenomenology and bridge the gap between particle and nuclear theory. For instance, numerical in-medium DGLAP evolution was validated on the example of HERMES measurements [398] and used to make the first prediction of the centrality dependence of light and heavy meson production in $e+A$ collisions at the EIC [399].
While the fully numerical approaches in the literature give a good description of jet quenching data, they do not provide insight into the types of logarithms that are being resummed or the fixed-order contributions to the modification of the measured cross sections. Most recently, a renormalization group (RG) analysis of collinear hadron production in nuclear DIS was performed in the limit where the parent parton energy $E$ is large while the medium opacity $L/\lambda$ remains small. This resulted in the derivation of a novel set of evolution equations [400] that resum leading $\ln(E/\mu^{2}L)$ enhanced medium contributions arising from multiple emissions near the endpoints of the splitting functions at first order in opacity. The RG evolution framework can be applied to study fragmentation in $e+A$ reactions as a function of $z_{h}=E_{h}/\nu$, as shown in the top panel of Fig. 15. The nuclear modification ratios, normalized by the inclusive cross sections to eliminate initial-state effects,
@@ IV.3 A renormalization group approach to energy loss
Figure 15: Top: Modification of the differential hadron cross section in SIDIS for several nuclei with different center-of-mass energies and vitualities [400]. Bottom: Ratio of $p_{T}$-differential DY production cross section on heavy to light nuclei compared to TMD models including in-medium broadening and radiative corrections [401].
A major hadron physics goal of the past decade has been to understand the three-dimensional structure of nucleons. Such structure is encoded in transverse momentum dependent parton distribution functions (TMDs) that can be extracted trough global analyses  [402, 403, 404, 405, 406]. Despite the progress in our understanding of TMD observables in elementary reactions with free nucleons, investigation of the three-dimensional structure of nuclei has just begun [407, 408].
Importantly, the first TMD investigation of the perturbative dynamics arising from forward scattering between the partonic constituents of the proton and the nuclear medium has recently appeared. The new contributions to low $p_{T}$ Drell-Yan production were obtained up to NLO in $\alpha_{s}$ and to first order in the medium opacity. The derivation includes both collisional broadening and initial-state energy loss. The collinear and rapidity divergences related to parton showers in the nucleus and the coherent nature of radiation due to the LPM effect lead to an in-medium RG equation that encodes the $p_{T}$ dependence of parton energy loss. In addition, the rapidity evolution leads to a Balitsky-Fadin-Kuraev-Lipatov evolution equation with restricted phase space for the forward scattering cross section.

### v2

@@ IV Radiative parton energy loss
Reactions with nuclei are essential for the study of QCD. Cold nuclear matter effects can be investigated in $e+A$ [24] and $p+A$ collisions [371, 372]. The modification of hadron production cross sections at low and intermediate $p_{T}$, known as the Cronin effect [37], was established as early as the mid 1970s. In the 1990s, fixed-target measurements of Drell-Yan and charmonia at Fermilab (see Sec. II) found a substantial suppression of the cross sections in nuclei at large $x_{F}$. While initially attributed only to nPDF effects, it was noted early on that attenuation of the production rates of such states is also compatible with CNM energy loss [373].
In the late 1990s and early 2000s important developments in many-body perturbative QCD produced theoretical frameworks to evaluate non-Abelian parton energy loss due to soft gluon emission [374, 375, 19, 10, 376, 377, 378, 379]. Subsequently, the energy loss mechanism has been extensively employed to study Drell-Yan [380, 41, 381, 42] and $J/\psi$ [373, 382, 13, 383] suppression at large $x_{F}$ in minimum bias $p+A$ collisions as well as in jet studies [384] in central $p+A$ collisions at very high energies.
In the remainder of this section, we describe the various energy loss regimes and observables that are sensitive to these regimes. We further touch upon recent progress on understanding in-medium radiation beyond the soft gluon emission limit and its connection to other perturbative approaches such as saturation.
@@ IV.1 Energy loss regimes
These processes are widely studied as one of the dominant mechanisms for high $p_{T}$ hadron and jet suppression in heavy-ion collisions. However, they may also play a significant role in forward particle production of quarkonium [382, 13, 383], light hadrons [385, 386], and heavy flavor mesons [233] in $p+A$ collisions.
@@
The medium-induced gluon radiation spectrum, the difference between gluon radiation in protons and nuclei, exhibits different scaling properties depending on the typical gluon formation time $t_{f}\sim\omega/k_{T}^{2}$, where $\omega$ and $k_{T}$ are the energy and the transverse momentum of the radiated gluon, respectively. This gluon formation time needs to be compared to the two other length scales appearing in the process, the parton mean free path in the medium, $\lambda$, and the length of the medium traversed, $L$. In practice, qualitative analytic behavior can be derived in the limits of very large parton energy and large numbers of scatterings. Numerical corrections can be large for lower energies and a smaller number of scatterings [10, 387].
It is useful to recall the different regimes identified in the medium-induced gluon spectrum depending on the gluon formation time $t_{f}$ [388]:
Incoherent or Bethe-Heitler regime
@@
In this regime, $t_{f}\lesssim\lambda$, each scattering center acts as an independent radiation source. This regime is important for very soft gluons, $\omega\ll\mu^{2}\lambda=\hat{q}\lambda^{2}$, where $\hat{q}=2\mu^{2}/\lambda$ is the medium transport coefficient and $\mu$ is the inverse of the interaction range [388, 389];
Partially coherent or LPM regime
@@
Here $\lambda\ll t_{f}\lesssim L$, and a group of scattering centers with $t_{f}/\lambda\gg 1$ acts as a single radiator, leading to relative suppression of the gluon radiation intensity with respect to the intensity in the incoherent regime. This condition leads to $\mu^{2}\lambda\ll\omega\lesssim\hat{q}L^{2}$. The average LPM energy loss scales parametrically as $\Delta E_{{}_{\textnormal{LPM}}}\propto\hat{q}L^{2}\ln(E/\mu^{2}L)$, with a logarithmic dependence on the parton energy;
Fully coherent regime
@@
Here $t_{f}\gg L$ and all scattering centers in the medium act as a coherent radiation source. In this large gluon energy regime, $\omega\gg\hat{q}L^{2}$, the average energy loss $\Delta E_{{}_{\textnormal{FCEL}}}\propto(\hat{q}\,L)^{1/2}\,(E/M_{T})$ becomes proportional to the parton energy $E$ [12, 13], overwhelming the LPM energy loss, $\Delta E_{{}_{\textnormal{FCEL}}}\gg\Delta E_{{}_{\textnormal{LPM}}}$. The phenomenology of FCEL has been investigated in quarkonium [382, 13, 383], light hadron [385, 386], and heavy flavor meson [233] production in $p+A$ collisions. The effects of FCEL, which explicitly break QCD factorization, are sizable and therefore cannot be neglected. Given this, including measurements of hadron production in $p+A$ collisions in global analyses of nPDFs [225, 224, 226] without proper implementation of FCEL might lead to biased nPDF extractions.
Bertsch-Gunion (BG)-like regime
@@
Initial-state interactions can partially cancel radiation effects, especially at low frequencies. In this regime, This medium-induced radiation [379] exhibits Bertsch-Gunion [390, 391] scaling properties, $\omega(dI/d\omega)\sim L$, without additional power suppressions. The type of energy loss is manifested in Drell-Yan, hadron, and jet suppression at large $x_{F}$ [41, 392, 393].
@@
Figure 12: Schematic log-log plot showing the parametric dependence of the $\alpha_{s}$-normalized induced radiation spectrum $(\omega/\alpha_{s})(dI/d\omega)$ of parton production in the color representation $R$ at small angles in the regimes of gluons formation times described here. Taken from Ref. [394].
@@
Figure 12 schematically summarizes the parametric dependence of the medium-induced radiation spectrum. It can be shown that if a parton suddenly decelerated (accelerated), leading to initial-state, IS, or (final-state, FS) energy loss, the medium-induced gluon spectrum decreases as $\omega(dI/d\omega)\sim(1/\omega)$ at $\omega\gg\hat{q}\,L^{2}$, leading to $\Delta E_{{}_{\textnormal{IS}}}=\Delta E_{{}_{\textnormal{FS}}}=\Delta E_{{}_{\textnormal{LPM}}}$. On the contrary, whenever the hard process appears to be the result of small-angle deflection of a high energy parton, as is typically the case for hadron production in $p+A$ collisions when both the initial-state and the final-state partons carry a color charge (see Fig. 13), the associated medium-induced gluon radiation arises from the interference of gluon radiation before and after the hard scattering. In that case, the induced spectrum is dominated by the fully coherent radiation regime, $t_{\textnormal{f}}\gg L$, leading to $\Delta E=\Delta E_{{}_{\textnormal{FCEL}}}$.
@@
Figure 13: Induced gluon radiation from a fast parton experiencing multiple scatterings in a dense nucleus. Taken from [394].
@@
Disentangling the different regimes of radiative parton energy loss is key to understanding its effects. At high collision energies, typically at the LHC, the effect of LPM energy loss in nuclear matter should be negligible because the incoming parton energy $E$ is extremely large in the nuclear rest frame, leading to negligible fractional energy loss, $\Delta E_{{}_{\rm LPM}}/E\sim\ln E/E\ll 1$. Thus the LPM regime would best be probed in low energy nuclear collisions.
@@
Initial state LPM $\hat{q}\,L^{2}\,\ln E$ Drell-Yan $h+A$ E906, COMPASS [42]
Final state LPM $\hat{q}\,L^{2}\,\ln E$ $h$, jets $e+A$ CLAS, HERMES, EIC [395]
FCEL $(\hat{q}\,L)^{1/2}\,(E/M_{T})$ $h$, jets $h+A$ SPS, FNAL, RHIC, LHC [382, 13, 383]
@@
Table 6: Energy loss processes in collisions on nuclear targets, their parametric dependencies, and the systems and facilities in which they might be probed. (No example of Bethe-Heitler observables is given.)
@@
IV.2 Cold nuclear matter transport coefficient $\hat{q}$
The strength of induced gluon radiation in cold nuclear matter is governed by the transport coefficient $\hat{q}$ within the Baier–Dokshitzer–Mueller–Peigné–Schiff and Zakharov (BDMPS–Z) formalism [^7] Note that in Eq. (5), $\tilde{\mu}^{2}$ already accounts for the typical transverse momentum squared.  [388, 19, 396, 389], which quantifies the scattering power of the nuclear medium. It is related to the gluon distribution $xg(x,Q^{2})$ inside each nucleon,
$\hat{q}\equiv\frac{\tilde{\mu}^{2}}{\lambda}=\frac{4\pi^{2}\,\alpha_{s}\,C_{R}}{N_{c}^{2}-1}\;\rho\,xg\bigl{(}x,Q^{2}\bigr{)}\,,$ (5)
where $Q^{2}\sim\Delta p_{T}^{2}$ denotes the characteristic transverse momentum scale, $C_{R}$ is the parton color factor ($C_{F}=(N_{c}^{2}-1)/2N_{c}$ for a quark and $C_{A}=N_{c}$ for a gluon), and $\rho$ is the nuclear density.
@@ IV.2 Cold nuclear matter transport coefficient $\hat{q}$
[…]
@@
This value is indirectly related to multiple parton scattering in the nuclear medium, $\Delta p_{T}^{2}\propto\hat{q}_{A}L_{A}-\hat{q}_{p}L_{p}$. As shown in Fig. 14, using world data on $J/\psi$ and Drell-Yan transverse momentum broadening in $h+A$ collisions from SPS to the LHC, the transport coefficient is found to be $\hat{q}_{0}=0.051\,\text{GeV}^{2}/\text{fm}$ or $\hat{q}_{0}=0.075\,\text{GeV}^{2}/\text{fm}$ depending on the assumed color state of the produced $c\overline{c}$ pair that fragments into a $J/\psi$ [15]. The lower value of $\hat{q}^{2}_{0}$ assumes color octet production while the larger value assumes that the pair is produced as a color singlet.
@@
IV.3 Renormalization group approach to energy loss
In the past decades, successful energy loss phenomenology was developed for both hot and cold QCD matter. Such theoretical models, however, have clear limitations. They cannot be systematically improved to higher perturbative orders and do not give a correct description of in-medium parton showers. For example, in the case of open heavy flavor production, the $g\rightarrow Q\bar{Q}$ process, which is especially important at high energies, is not included. To address these deficiencies, full medium-induced splitting kernels [398, 399] were derived using effective field theory and lightcone wavefunction approaches. These analogs of DGLAP splitting functions in matter further require implementation in cross section evaluations that go beyond energy loss phenomenology and bridge the gap between particle and nuclear theory. For instance, numerical in-medium DGLAP evolution was validated on the example of HERMES measurements [400] and used to make the first prediction of the centrality dependence of light and heavy meson production in $e+A$ collisions at the EIC [401].
While the fully numerical approaches in the literature give a good description of jet quenching data, they do not provide insight into the types of logarithms that are being resummed or the fixed-order contributions to the modification of the measured cross sections. Most recently, a renormalization group (RG) analysis of collinear hadron production in nuclear DIS was performed in the limit where the parent parton energy $E$ is large while the medium opacity $L/\lambda$ remains small. This resulted in the derivation of a novel set of evolution equations [402] that resum leading $\ln(E/\mu^{2}L)$ enhanced medium contributions arising from multiple emissions near the endpoints of the splitting functions at first order in opacity. The RG evolution framework can be applied to study fragmentation in $e+A$ collisions as a function of $z_{h}=E_{h}/\nu$, as shown in the top panel of Fig. 15. The nuclear modification factors, normalized by the inclusive cross sections to eliminate initial-state effects,
@@ IV.3 Renormalization group approach to energy loss
Figure 15: Top: Modification of the differential hadron cross section in SIDIS for several nuclei with different center-of-mass energies and vitualities [402]. Bottom: Ratio of $p_{T}$-differential Drell-Yan production cross sections on heavy to light nuclei compared to TMD models including in-medium broadening and radiative corrections [403].
A major hadronic physics goal of the past decade has been to understand the three-dimensional structure of nucleons. Such structure is encoded in transverse momentum dependent parton distribution functions (TMDs) that can be extracted through global analyses  [404, 405, 406, 407, 408]. Despite the progress in our understanding of TMD observables in elementary reactions with free nucleons, investigation of the three-dimensional structure of nuclei has just begun [409, 410].
Importantly, the first TMD investigation of the perturbative dynamics arising from forward scattering between the partonic constituents of the proton and the nuclear medium has recently appeared. The new contributions to low $p_{T}$ Drell-Yan production were obtained up to NLO in $\alpha_{s}$ and to first order in the medium opacity. The derivation includes both collisional broadening and initial-state energy loss. The collinear and rapidity divergences related to parton showers in the nucleus and the coherent nature of radiation due to the LPM effect lead to an in-medium RG equation that encodes the $p_{T}$ dependence of parton energy loss. In addition, the rapidity evolution leads to a BFKL evolution equation with restricted phase space for the forward scattering cross section.